# A Longitudinal Study on the Attitudes of Gay Men in Beijing Towards Gay Social Media Platforms: Lonely Souls in the Digital Concrete Jungle

### v2

@@ 4.1.2. Advances in technology, and the rise of tensions and critique
In addition, age- and education-based discrimination was common. Eleven participants reported experiences of education shaming (based on academic credentials). Five participants reported age shaming. For example:
@@ 4.1.3. Pragmatism and Development
Table 4. A summary of social applications used by participants across four years (2014, 2017, 2020, and 2023). The symbol ‘—‘ indicates that no specific application was mentioned or used by the participant for that year.
Participant 2014 2017 2020 2023
1 Aloha, Blued, Finka Ginder
2 Aloha, Blued, Finka Aloha, Blued Blued, Finka Wechat, Tinder
3 Aloha, Blued, Finka Blued
4 Aloha, Finka Aloha, Finka Blued, Aloha Hinge, Blued
5 Aloha, Blued, Finka Aloha, Blued Blued Hinge
6 Aloha, Blued, Finka Aloha, Blued, Finak Finka QSmatch
7 Aloha, Blued, Finka Aloha, Blued Blued, Finka QSmatch
8 Aloha, Finka Aloha, Finka Blued Blued
9 Aloha, Blued, Finka Blued Aloha, Finka Finka
10 Aloha, Finka Aloha, Finka Aloha, Finka QSmatch
11 Aloha, Blued, Finka Blued, Zank Aloha, Blued Blued
12 Aloha, Blued, Finka Blued, Finka Aloha, Finka
13 Aloha, Blued Aloha, Blued Aloha, Finka Blued
14 Aloha, Blued Aloha, Blued Aloha, Finka Blued
15 Blued, Zank Blued, Zank Aloha, Finka, Tinder Tinder, QSmatch
16 Blued Blued, Finka Blued, Aloha
17 Aloha, Finka Aloha, Finka
@@ 5. Discussion
This study presents a nine-year, mixed-method longitudinal investigation into the evolving attitudes of gay men in China toward gay social applications. Our findings reveal a clear trajectory: an early phase of optimistic adoption, a middle phase of ambivalence and critique, and a recent phase of fragmented, strategic use. This evolution not only reflects the development of the technological platforms themselves but also mirrors the complex interplay among China’s sociocultural context, users’ psychological needs, and digital technologies. In the following sections, we discuss the findings of our study from multiple perspectives.

### v3

@@ 4.2.2. Advances in technology, and the rise of tensions and critique
In addition, age-based and education-based discrimination was common. Eleven participants reported experiences of education shaming (based on academic credentials). Five participants reported age shaming. For example:
@@ 5. Discussion
Our primary explanatory lens centers on users’ adaptive portfolio strategies in response to evolving sociocultural pressures, with platform affordances serving as contextual triggers rather than the sole drivers of attitudinal change. This study presents a nine-year, mixed-method longitudinal investigation into the evolving attitudes of gay men in China toward gay social applications. Our findings reveal a clear trajectory: an early phase of optimistic adoption, a middle phase of ambivalence and critique, and a recent phase of fragmented, strategic use. This evolution not only reflects the development of the technological platforms themselves but also mirrors the complex interplay among China’s sociocultural context, users’ psychological needs, and digital technologies. In the following sections, we discuss the findings of our study from multiple perspectives.
@@ 7. Conclusion
Appendix A Researcher Positionality Statement
This study adheres to the principles of objectivity, neutrality, and ethical responsibility in academic research. It aims to understand and document the evolving attitudes of men who have sex with men (MSM) in China toward queer social platforms through rigorous and scientifically grounded methodologies. The researchers recognize that topics related to sexual minority groups may involve complex social sensitivities and policy contexts within the Chinese sociocultural environment.
  1. (1)
    This study adopts perspectives from Human–Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW), with a focus on objectively examining technology use behaviors, changes in user experience, and their interactions with the sociotechnical environment. The research aims to understand user practices and attitudinal shifts, as well as their underlying technological and social drivers, thereby generating insights to inform the design of digital environments that are more inclusive, supportive, and care-oriented.
  2. (2)
    We fully respect the dignity, autonomy, and privacy of all research participants. The objective of this study is to understand—rather than evaluate—the digital practices of MSM in China. We acknowledge the unique challenges that this community may encounter in everyday life, and we commit to contributing to knowledge through rigorous and ethical research practices while avoiding any action that could reinforce stigmatization or cause harm.
  3. (3)
    This study has been approved by the Institutional Review Board (IRB) of the affiliated institution (name withheld for anonymity) and strictly complies with core ethical standards, including informed consent, anonymization, data confidentiality, and the right of participants to withdraw at any time without justification. All interviewers received professional training to minimize any potential risks associated with participation.
  4. (4)
    This study does not make value judgments or engage in advocacy regarding any political or social systems, cultural norms, or policies. It strictly complies with Chinese laws and regulations as well as the policies of the affiliated academic institution. All data collection and analysis were conducted within the boundaries permitted by applicable legal and policy frameworks.
Appendix B Longitudinal Evolution of Higher-Order Themes and Sub-themes (2014–2023)
Table 5. Longitudinal Evolution of Higher-Order Themes and Sub-themes (2014–2023)
Higher-Order Theme Sub-theme Definition 2014 2017 2020 2023 Trend Sample Quote (Exemplar)
Technological Affordances Interface Utility Satisfaction with UI, speed, and features. 6 12 15 16 ↑ ($\chi^{2}=14.8$) “The filter function is now incredibly precise.” (P3, 2023)
Privacy Control Mechanisms to hide distance, photos, or tracks. 4 9 13 15 ↑ ($\chi^{2}=16.3$) “I can finally burn photos after viewing.” (P11, 2020)
Social Ecology Community Belonging Feeling of being connected to a collective. 15 16 8 3 ↓ ($\chi^{2}=26.5$) “It felt like a digital home back then.” (P1, 2014)
Deception & Risk Encountering fake profiles or scams. 2 5 12 14 ↑ ($\chi^{2}=21.4$) “Half the profiles are bots or scammers.” (P9, 2023)
Commodification Interactions driven by transactions/money. 0 3 9 13 ↑ ($\chi^{2}=24.1$) “Everything requires a VIP subscription now.” (P5, 2023)
Identity Negotiation Liberation Freedom to express sexual orientation. 14 15 10 6 ↓ ($\chi^{2}=12.9$) “I could finally be myself.” (P8, 2017)
Algorithmic Fatigue Tiredness from standardized presentation. 0 2 8 15 ↑ ($\chi^{2}=29.7$) “I am tired of being calculated by the system.” (P7, 2023)
Note: Values represent the number of participants ($N=17$) mentioning each sub-theme. Arrows indicate directional change; $\chi^{2}$ values are provided for descriptive support only.
Appendix C Quantitative Analysis Results
To rigorously examine the evolutionary trajectory of user attitudes and address concerns regarding the descriptive nature of survey data, we implemented a multi-layered quantitative framework comprising reliability and validity verification, Linear Mixed Models (LMM), and effect size calculation. First, we established the psychometric quality of our instrument through Exploratory Factor Analysis (EFA), which confirmed a robust two-factor structure distinguishing “Technological Experience” from “Social Outcomes.” Internal consistency for these factors was high, with Cronbach’s $\alpha$ coefficients of 0.89 for the technological sub-scale and 0.83 for the social sub-scale, indicating strong reliability. Furthermore, to verify construct validity, we assessed the relationship between the specific item ratings and the global “Overall Evaluation” (Q11). A Pearson correlation analysis revealed a significant positive correlation ($r=0.65,p<.001$) between the 10-item composite mean and the overall evaluation score, confirming that the specific dimensions measured are statistically valid predictors of the users’ general sentiment toward these platforms.
To test the core hypothesis that technical optimization has not translated into social well-being, we employed Linear Mixed Models (LMM) to control for within-subject correlations and missing data inherent in longitudinal designs. The analysis revealed a highly significant interaction effect between Year and Dimension Type ($p<.001$). While scores for Technological Experience followed a linear upward trajectory, Social Outcomes exhibited a distinct inverted U-shape that peaked in 2017 before declining. Specifically, the interaction term for 2023 indicated that the trajectory for technical experience significantly diverged from social outcomes, with technical scores increasing by a net magnitude of approximately 3.65 points relative to the social baseline trend ($\beta=5.57,p<.001$). This statistical divergence provides empirical support for the “Function-Social Gap,” demonstrating that the disparity between improved app usability and declining community sense is a systematic phenomenon rather than random fluctuation.
The analysis of the “Overall Evaluation” (Q11) further corroborated this trajectory of disenchantment. Consistent with the decline in specific social dimensions, the LMM results for the overall evaluation indicated a significant negative shift over the decade. Relative to the 2014 baseline ($M=6.47$), the overall user sentiment in 2023 dropped significantly ($\beta=-2.91,p<.001$), reaching a low of $M=3.54$. Notably, the model showed no significant difference between 2014 and 2020 ($\beta=0.03,p=.92$), suggesting that while the erosion of social value began earlier, the critical deterioration in overall sentiment accelerated precipitously in the post-2020 era.
Finally, we quantified the practical significance of these shifts using Cohen’s $d$ to measure the magnitude of change between the 2014 baseline and the 2023 final wave. The results highlighted a dramatic polarization in user experience. Technological dimensions showed large positive effects, with satisfaction regarding Functionality ($d=2.54$) and Interface Design ($d=2.33$) improving by over two standard deviations. Conversely, social dimensions suffered severe deterioration, most notably in the perceived feasibility of Long-term Relationships (d=2.33d=-2.33) and the sense of Belonging (d=2.30d=-2.30). The Overall Evaluation similarly reflected a substantial negative effect size of d=1.82d=-1.82. Collectively, these inferential statistics portray a robust picture of “Technical Success but Social Failure,” statistically validating the qualitative observation that these platforms have succeeded as functional tools while increasingly failing as social infrastructures.
Table 6. Results of Linear Mixed Effects Model (LMM) Analysis
Parameter Estimate ($\beta$) Std. Err. z-value p-value 95% CI
Model 1: Interaction Analysis
Intercept 6.275 0.168 37.33 <.001<.001 5.95 6.60
Year [2023] (Main Effect) -1.916 0.234 -8.19 <.001<.001 -2.37 -1.46
Year [2023] $\times$ Type [Tech] 5.574 0.328 16.97 <.001<.001 4.93 6.22
Model 2: Overall Evaluation (Q11)
Intercept 6.471 0.245 26.43 <.001<.001 5.99 6.95
Year [2017] 1.71 0.337 5.07 <.001<.001 1.05 2.37
Year [2020] 0.033 0.331 0.10 .92 -0.62 0.68
Year [2023] -2.911 0.353 -8.25 <.001<.001 -3.60 -2.22
Table 7. Descriptive Statistics, Reliability, and Effect Sizes by Dimension
Factor / Dimension Reliability 2014 Mean (SD) 2023 Mean (SD) Change Cohen’s $d$
Technological Experience 0.89 5.32 (0.97) 9.08 (0.42) +3.76 +2.59
Concise Interface 6.29 (1.05) 9.08 (0.49) +2.79 +2.33
Privacy Protection 5.18 (1.78) 9.31 (0.48) +4.13 +1.87
Functionality 4.82 (1.19) 8.85 (0.55) +4.03 +2.54
Social Outcomes 0.83 6.44 (0.88) 4.31 (0.89) -2.13 -2.05
Belonging 6.29 (1.21) 2.46 (1.27) -3.83 -2.30
Long-term Relationship 4.18 (1.38) 1.00 (1.08) -3.18 -2.33
Social Interaction 8.53 (1.28) 5.85 (1.21) -2.68 -2.45
Enjoyment 7.06 (1.34) 5.31 (0.48) -1.75 -1.22
Mental Health 6.59 (0.94) 5.92 (0.86) -0.67 -0.39
Self-acceptance 5.00 (0.71) 5.54 (1.05) +0.54 +0.56
Overall Evaluation (Q11) N/A 6.47 (1.07) 3.54 (1.05) -2.93 -1.82